### v1

# New Taipei City, 251301, Taiwan

Abstract
Children often suffer wrist trauma in daily life, while they usually need radiologists to analyze and interpret X-ray images before surgical treatment by surgeons. The development of deep learning has enabled neural networks to serve as computer-assisted diagnosis (CAD) tools to help doctors and experts in medical image diagnostics. Since the You Only Look Once Version-8 (YOLOv8) model has obtained the satisfactory success in object detection tasks, it has been applied to various fracture detection. This work introduces four variants of Feature Contexts Excitation-YOLOv8 (FCE-YOLOv8) model, each incorporating a different FCE module (i.e., modules of Squeeze-and-Excitation (SE), Global Context (GC), Gather-Excite (GE), and Gaussian Context Transformer (GCT)) to enhance the model performance. Experimental results on GRAZPEDWRI-DX dataset demonstrate that our proposed YOLOv8+GC-M3 model improves the mAP@50 value from 65.78% to 66.32%, outperforming the state-of-the-art (SOTA) model while reducing inference time. Furthermore, our proposed YOLOv8+SE-M3 model achieves the highest mAP@50 value of 67.07%, exceeding the SOTA performance. The implementation of this work is available at https://github.com/RuiyangJu/FCE-YOLOv8.
Keywords:
Deep Learning Computer Vision Object Detection Fracture Detection Medical Image Processing Medical Image Diagnostics You Only Look Once (YOLO) Feature Contexts Excitation
1 Intorduction
Wrist trauma is common among children [14, 35]. Meena et al. [32] have shown that most fractures occur around 2 centimeters from the distal radius near the joint. If they are not treated promptly and effectively, such trauma will result in wrist joint deformities, limited joint motion, and chronic pain [1]. In severe cases, an incorrect diagnosis would lead to lifelong complications and inconvenience [24].
AI models play a crucial role in the medical field [30]. With the development of deep learning, neural networks are increasingly used as computer-assisted diagnosis (CAD) tools [25, 31, 39, 40] to aid doctors and experts in analyzing medical images (e.g., X-ray images) [47]. Object detection networks can accurately predict fractures and reduce the probability of misdiagnosis. Although two-stage object detection networks [9, 10, 37] obtain the excellent performance, one-stage object detection networks [36] offer faster inference time. With the continued release of You Only Look Once (YOLO) series models, YOLOv8 [19] and YOLOv9 [43] models can perform object detection efficiently on low-computing platforms [22].
Since the release of the GRAZPEDWRI-DX [33] dataset by the Medical University of Graz, Ju et al. [21] have been the first to deploy YOLOv8 models to predict pediatric wrist fractures. This deployment helped surgeons to interpret fractures in X-ray images, reducing misdiagnosis and providing a better information base for the surgery. Chien et al. [5] developed YOLOv8-AM models by incorporating different attention modules [29, 44, 46, 48] into YOLOv8 [19], which improved the model performance in fracture detection. In addition, Chien et al. [4] applied YOLOv9 to this dataset, achieving the state-of-the-art (SOTA) model performance. However, there remains a great possibility for improving the performance of current SOTA models.
Models Params FLOPs mAPval 50
Inference
Time
YOLOv8+ResCBAM(SOTA) 53.87M 196.2G 65.78% 8.7ms
YOLOv8+GC-M3(Ours) 43.85M 165.6G 66.32% 7.9ms
YOLOv8+GCT-M2(Ours) 43.64M 165.4G 65.67% 15.3ms
YOLOv8+SE-M3(Ours) 43.85M 165.6G 67.07% 15.3ms
YOLOv8+GE-M2(Ours) 43.65M 165.4G 65.99% 15.3ms
Figure 1: GRAZPEDWRI-DX Dataset mAP@50 vs. Inference Time. The input image size is 1024; the sizes of YOLOv8 and its improved models are all large (L), and the size of YOLOv9 model is extended (E). The inference time per image is measured on NVIDIA GeForce RTX 3090 GPUs. Our proposed YOLOv8+GE, YOLOv8+GC, and YOLOv8+SE models outperform the current SOTA model on mAP@50 value.
Studies [3, 16, 17, 38] have shown that incorporating different Feature Contexts Excitation (FCE) modules into Convolutional Neural Networks (CNNs) can enhance the model performance. For instance, the Squeeze-and-Excitation (SE) [17] module can reduce the Top-5 error from 7.48% to 6.62% for the ResNet-50 [13] model on the ImageNet dataset [6], with only a slight increase in computational cost. The Global Context (GC) [3] module can enhance ResNet-50 [13] model performance by 2.7% $\uparrow$ on $AP^{bbox}$ and 2.4% $\uparrow$ on $AP^{mask}$ for object detection on the COCO 2017 [27] validation set. The ResNet-50 [13] model with the Gather-Excite (GE) [16] module outperforms its 101-layer counterpart on the ImageNet dataset [6] without adding any learnable parameters. Incorporating the Gaussian Context Transformer (GCT) [38] module into Mask R-CNN [12] with ResNet-50 [13] model increases $AP^{bbox}_{0.5}$ from 38.2% to 39.4% on the COCO 2017 [27] validation set. Therefore, we propose FCE-YOLOv8 models for pediatric wrist fracture detection, each incorporating SE [17], GC [3], GE [16], and GCT [38] modules into YOLOv8, respectively, to enhance the model performance, and the results are shown in Fig. 1. This paper makes the following contributions:
  • (1)
    Introduces four different FCE modules to YOLOv8 model architecture, respectively, and proposes the FCE-YOLOv8 model, specifically designed for pediatric wrist fracture detection. One variant of our proposed FCE-YOLOv8 models, YOLOv8+SE, obtains mAP@50 values surpassing the SOTA level on the GRAZPEDWRI-DX public dataset.
  • (2)
    Since the main improvements in YOLOv8 can be categorized into three methods, we conduct the experiment to evaluate the most efficient method for each variant of our proposed FCE-YOLOv8 models that incorporates different FCE modules.
  • (3)
    Our proposed YOLO+GC model achieves the second-best performance on mAP@50 value on the GRAZPEDWRI-DX public dataset with only a slight increase in inference time, suitable as CAD tools to assist doctors and experts in interpreting pediatric wrist trauma X-ray images in real-world.
The rest of this paper is organized as follows: Section 2 reviews various research on pediatric wrist fracture detection based on YOLO models and related modules. Section 3 introduces FCE-YOLOv8 model architectures on three different improved methods, and FCE modules used in this work. Section 4 presents the experiment conducted to evaluate the most efficient improved method for different FCE-YOLOv8 models, and compares our proposed models with other SOTA models. Section 5 discusses the limitations of this work, particularly the class imbalance in the used dataset. Finally, Section 6 summarises the outcomes of this paper and suggests directions for future work.
2 Related Work
2.1 YOLO Models for Pediatric Wrist Fracture Detection
Pediatric wrist fracture detection is an important topic in medical image diagnostic tasks. Publicly available datasets [34] for wrist fracture detection in adults are limited, and even fewer datasets [11] for pediatrics. Nagy et al [33] published a pediatric related dataset named the GRAZPEDWRI-DX dataset. Hržić et al. [15] demonstrated that YOLOv4 [2] can enhance the accuracy of pediatric wrist trauma diagnoses in X-ray images. Ju et al. [21] developed an application employing YOLOv8 to assist doctors in interpreting pediatric wrist tramua X-ray images. Chien et al. [5] introduced four attention modules, including Convolutional Block Attention Module (CBAM) [46], Efficient Channel Attention (ECA) [44], Global Attention Mechanism (GAM) [29], and Shuffle Attention (SA) [48] to improve the model performance, where YOLOv8+ResCBAM achieving the SOTA performance on the mAP@50 value. In addition, Chien et al. [4] applied YOLOv9 [43] to this dataset, and achieved excellent performance.
2.2 Related Modules
FCE modules (i.e., SE [17], GC [3], GE [16], and GCT [38]) used in this work can enhance the feature modeling capabilities of neural networks in object detection tasks. (1) SE [17], introduced by Hu et al. in 2017, can improve the network ability to select channel features. The main contribution of this module is to dynamically adjust the channel weights of the feature map through the “Squeeze” and “Excitation” mechanisms. (2) GC [3], proposed by Cao et al. in 2019, can enhance the convolutional operations by efficiently capturing the global context information, which is the simplified and optimized version of Non-local (NL) [45] block. (3) GE [16], presented by Hu et al. in 2018, improves the SE module to better integrate the spatial and channel information. Compared to SE, GE captures more detailed spatial information and enables spatial adjustments at more granularity. (4) GCT [38], introduced by Ruan et al. in 2021, incorporates the Gaussian distribution into attention mechanism to capture the global context. This module uses the kernel function of Gaussian distribution to model the global context correlations, which further improves the accuracy of context capturing.
Figure 2: Network architecture of improved method-1 (M1): adding one FCE module to the Backbone component of YOLOv8.
Figure 3: Network architecture of improved method-2 (M2): adding one FCE module to the Head component of YOLOv8.
Figure 4: Network architecture of improved method-3 (M3): adding four FCE modules to the Head component of YOLOv8.
Figure 5: Detailed illustration of FCE modules used in this work: (a) SE module, (b) GC module, (c) GE module, and (d) GCT module.
3 Proposed Method
3.1 Network Architecture
The network architecture of YOLOv8 [19] comprises three components: Backbone, Neck, and Head. The Backbone component employs the Cross Stage Partial (CSP) [42] strategy, which divides the feature map into two parts. One part uses convolution operations, while the other is concatenated with the convolutional output of the first part. In the Neck component, YOLOv8 applies a combination of Feature Pyramid Network (FPN) [26] and Path Aggregation Network (PAN) [28] for multi-scale feature fusion. This component deletes convolution operations during the upsampling stage, allowing the top feature maps to retain more information by utilizing more layers, while minimizing the loss of location information in the bottom feature maps due to fewer convolutional layers. In addition, YOLOv8 employs a decoupled Head [8] to separate the classification and detection tasks. It replaces the anchor-based method used in YOLOv5 [20] with an anchor-free method [7], which locates objects by recognizing their center points and predicting the distance from the center to the bounding box.
To enhance global feature extraction and capture comprehensive information from medical images, we have designed three improved methods by incorporating different FCE modules into different components of YOLOv8 network architecture, as illustrated in Figs. 2, 3, and 4, respectively. Specifically, in the improved method 1 (M1), FCE modules (i.e., SE, GC, GE, and GCT) are added after the Spatial Pyramid Pooling - Fast (SPPF) [19] layer in the Backbone component. In the improved method 2 (M2), different FCE modules are integrated into the final C2f [19] module of the Head component. While in the improved method 3 (M3), FCE modules are added after each of the four C2f modules in the Head component. For each FCE-YOLOv8 model, we select the most appropriate method based on experimental results, as discussed in Section 4.5.
3.2 Feature Contexts Excitation Modules
This work uses four different FCE modules, including SE [17], GC [3], GE [16], and GCT [38].
SE enhances the representational capacity of neural networks, including YOLOv8, by dynamically adjusting the weights of feature channels based on their importance. This enables YOLOv8 to focus more on relevant channel features while disregarding less significant ones. Specifically, SE employs the “Squeeze” and ”Excitation” operations to enhance the representation of meaningful features, improving both classification and detection performance. In object detection tasks (e.g., fracture detection), integrating SE into YOLOv8 can significantly improve the model performance
GC is highly suitable for fracture detection tasks due to its flexible properties and ease of integration into different neural network architectures, including the YOLOv8 model. This module combines the advantages of the Simplified Non-Local (SNL) [3] block, which effectively models long-range dependencies, and the SE module, known for its computational efficiency. In addition, GC follows the NL [45] block and aggregates the global context across all locations, enabling the capture of long-range dependencies. Furthermore, GC is lightweight and can be applied to multiple layers, enhancing the model ability to capture long-range dependencies with only a slight increase in computational cost.
GE can efficiently aggregate the global context information while maintaining spatial resolution, without significantly increasing the model computational demands. This makes it possible for use in CAD tools on the low-computing-power platform for fracture prediction. By gathering the global context information from the entire feature map, GE captures broader context effectively. In addition, different from other methods [18, 41] that aggregate global context information, GE preserves spatial resolution, which is useful for tasks requiring precise localization, such as fracture detection. Overall, GE can enhance the feature representation capabilities of the network, improving YOLOv8 model performance.
GCT can enhance the ability of YOLOv8 to capture context information by incorporating the Gaussian filter. This integration improves the capacity of the model for data modeling and allows for more effective utilization of training data by weighting both local and global context information. In addition, this module increases the efficiency of data utilization, reducing the demands for extensive datasets. Furthermore, by replacing the traditional self-attention mechanism with the Gaussian filter, GCT decreases computational complexity and enhances efficiency, particularly with large-scale data. In general, the Gaussian filter contributes to the model stability during training, mitigating issues such as gradient explosion or vanishing gradients. This results in improved convergence speed and overall performance of YOLOv8.
Table 1: Experimental results of selecting improved methods for pediatric wrist fracture detection using the GRAZPEDWRI-DX dataset for YOLOv8+SE model. The best performance and the 2nd best performance are in red and blue colors, respectively.
Model Method
Model
Size
Input
Size
[…]
Inference
Time
YOLOv8+SE M1 Small 640 11.30M 29.1G 62.18% 2.6ms
Medium 640 26.12M 79.8G 62.35% 5.1ms
Large 640 43.93M 166.3G 62.93% 7.6ms
YOLOv8+SE M2 Small 640 11.17M 28.7G 61.45% 2.5ms
Medium 640 25.90M 79.1G 61.75% 5.0ms
Large 640 43.67M 165.5G 64.08% 8.0ms
YOLOv8+SE M3 Small 640 11.19M 28.7G 61.72% 2.6ms
Medium 640 25.95M 79.2G 62.63% 5.0ms
Large 640 43.85M 165.6G 62.79% 7.6ms
YOLOv8+SE M1 Small 1024 11.30M 29.1G 63.24% 4.9ms
Medium 1024 26.12M 79.8G 63.70% 10.1ms
Large 1024 43.93M 166.3G 64.55% 15.4ms
YOLOv8+SE M2 Small 1024 11.17M 28.7G 62.90% 5.0ms
Medium 1024 25.90M 79.1G 63.68% 10.0ms
Large 1024 43.67M 165.5G 64.40% 14.7ms
YOLOv8+SE M3 Small 1024 11.19M 28.7G 62.77% 5.1ms
Medium 1024 25.95M 79.2G 64.10% 10.0ms
Large 1024 43.85M 165.6G 67.07% 15.3ms
Table 2: Experimental results of selecting improved methods for pediatric wrist fracture detection using the GRAZPEDWRI-DX dataset for YOLOv8+GC model. The best performance and the 2nd best performance are in red and blue colors, respectively.
Model Method
Model
Size
Input
Size
[…]
Inference
Time
YOLOv8+GC M1 Small 640 11.34M 29.1G 60.68% 1.7ms
Medium 640 26.17M 79.9G 63.07% 2.5ms
Large 640 43.97M 166.3G 64.55% 3.5ms
YOLOv8+GC M2 Small 640 11.21M 28.7G 61.30% 1.9ms
Medium 640 25.95M 79.2G 62.71% 2.7ms
Large 640 43.70M 165.5G 64.54% 3.5ms
YOLOv8+GC M3 Small 640 11.24M 28.7G 62.19% 1.9ms
Medium 640 26.03M 79.2G 62.77% 2.7ms
Large 640 43.85M 165.6G 63.97% 3.6ms
YOLOv8+GC M1 Small 1024 11.34M 29.1G 63.24% 2.9ms
Medium 1024 26.17M 79.9G 63.87% 5.1ms
Large 1024 43.97M 166.3G 64.85% 7.8ms
YOLOv8+GC M2 Small 1024 11.21M 28.7G 63.51% 2.8ms
Medium 1024 25.95M 79.2G 65.25% 5.1ms
Large 1024 43.70M 165.5G 65.70% 7.8ms
YOLOv8+GC M3 Small 1024 11.24M 28.7G 64.16% 2.8ms
Medium 1024 26.03M 79.2G 64.33% 5.2ms
Large 1024 43.85M 165.6G 66.32% 7.9ms
Table 3: Experimental results of selecting improved methods for pediatric wrist fracture detection using the GRAZPEDWRI-DX dataset for YOLOv8+GE model. The best performance and the 2nd best performance are in red and blue colors, respectively.
Model Method
Model
Size
Input
Size
[…]
Inference
Time
YOLOv8+GE M1 Small 640 11.28M 29.1G 62.44% 2.6ms
Medium 640 26.10M 79.8G 63.00% 5.1ms
Large 640 43.91M 166.3G 64.15% 7.6ms
YOLOv8+GE M2 Small 640 11.15M 28.7G 62.07% 2.6ms
Medium 640 25.87M 79.1G 63.20% 5.1ms
Large 640 43.65M 165.4G 64.02% 7.6ms
YOLOv8+GE M3 Small 640 11.16M 28.7G 61.89% 2.7ms
Medium 640 25.90M 79.1G 62.22% 5.2ms
Large 640 43.68M 165.5G 62.42% 7.7ms
YOLOv8+GE M1 Small 1024 11.28M 29.1G 63.72% 5.2ms
Medium 1024 26.10M 79.8G 63.81% 10.0ms
Large 1024 43.91M 166.3G 64.34% 15.4ms
YOLOv8+GE M2 Small 1024 11.15M 28.7G 62.51% 5.1ms
Medium 1024 25.87M 79.1G 64.28% 10.0ms
Large 1024 43.65M 165.4G 65.99% 15.3ms
YOLOv8+GE M3 Small 1024 11.16M 28.7G 62.42% 4.8ms
Medium 1024 25.90M 79.1G 62.63% 10.3ms
Large 1024 43.68M 165.5G 64.35% 15.7ms
Table 4: Experimental results of selecting improved methods for pediatric wrist fracture detection using the GRAZPEDWRI-DX dataset for YOLOv8+GCT model. The best performance and the 2nd best performance are in red and blue colors, respectively.
Model Method
Model
Size
Input
Size
[…]
Inference
Time
YOLOv8+GCT M1 Small 640 11.27M 29.1G 61.91% 2.7ms
Medium 640 26.08M 79.8G 62.64% 5.1ms
Large 640 43.90M 166.3G 63.70% 7.6ms
YOLOv8+GCT M2 Small 640 11.14M 28.7G 61.36% 2.6ms
Medium 640 25.86M 79.1G 62.19% 5.1ms
Large 640 43.64M 165.4G 63.42% 7.6ms
YOLOv8+GCT M3 Small 640 11.14M 28.7G 52.07% 2.7ms
Medium 640 25.86M 79.1G 53.60% 5.2ms
Large 640 43.64M 165.4G 57.88% 7.6ms
YOLOv8+GCT M1 Small 1024 11.27M 29.1G 62.49% 5.1ms
Medium 1024 26.08M 79.8G 63.47% 10.1ms
Large 1024 43.90M 166.3G 63.87% 15.3ms
YOLOv8+GCT M2 Small 1024 11.14M 28.7G 63.51% 5.0ms
Medium 1024 25.86M 79.1G 64.46% 10.1ms
Large 1024 43.64M 165.4G 65.67% 15.3ms
YOLOv8+GCT M3 Small 1024 11.14M 28.7G 52.53% 5.1ms
Medium 1024 25.86M 79.1G 53.70% 10.1ms
Large 1024 43.64M 165.4G 58.15% 15.7ms
Table 5: Quantitative comparison (Params/FLOPs/F1-Score/mAP@50/mAP@50-95/Inference time per image) with other SOTA models for pediatric wrist fracture detection on the GRAZPEDWRI-DX dataset. The sizes of YOLOv8 and its improved models are all large (L), and the size of YOLOv9 model is extended (E). The best performance and the 2nd best performance are in red and blue colors, respectively.
Model
Input
Size
[…]
Inference
Time
YOLOv8 1024 43.61M 164.9G 0.62 63.58% 7.7ms
YOLOv8+SA 1024 43.64M 165.4G 0.63 64.25% 8.0ms
YOLOv8+ECA 1024 43.64M 165.5G 0.65 64.24% 7.7ms
YOLOv8+GAM 1024 49.29M 183.5G 0.65 64.26% 12.7ms
YOLOv8+ResGAM 1024 49.29M 183.5G 0.64 64.98% 18.1ms
YOLOv8+ResCBAM 1024 53.87M 196.2G 0.64 65.78% 8.7ms
YOLOv9 1024 69.42M 244.9G 0.66 65.62% 16.1ms
YOLOv8+GC-M3 1024 43.85M 165.6G 0.66 66.32% 7.9ms
YOLOv8+GCT-M2 1024 43.64M 165.4G 0.64 65.67% 15.3ms
YOLOv8+SE-M3 1024 43.85M 165.6G 0.66 67.07% 15.3ms
YOLOv8+GE-M2 1024 43.65M 165.4G 0.64 65.99% 15.3ms
Figure 6: Visualization of the accuracy of predicting each class using our proposed FCE-YOLOv8 models and other SOTA models on the GRAZPEDWRI-DX dataset with the input image size of 1024: (a) YOLOv8+ResCBAM, (b) YOLOv9, (c) YOLOv8+GC, (d) YOLOv8+GCT, (e) YOLOv8+SE, and (f) YOLOv8+GE.
Figure 7: Examples of prediction results by our proposed models and other SOTA models, and Ground-Truth images of pediatric wrist fracture detection on X-ray Images.
4 Experiments
4.1 Datasets
GRAZPEDWRI-DX dataset, introduced by Nagy et al. [33], is a publicly available collection of pediatric wrist trauma X-ray images provided by the Medical University of Graz. This dataset was collected by radiologists at the University Hospital Graz between 2008 and 2018, and includes 20,327 X-ray images. In addition, it covers 6,091 pediatric patients and contains 10,643 studies, with a total of 74,459 image labels and 67,771 annotated objects. Nagy et al. [33] emphasize that prior to the release of their dataset, there are few publicly available relevant pediatric datasets. Therefore, we only used this dataset for our experiments.
4.2 Evaluation Metrics
Fracture detection on pediatric wrist X-ray images belongs to object detection tasks, and six evaluation metrics in object detection tasks are applicable, including the model parameters (Params), floating-point operations (FLOPs), F1-Score, mean average precision at 50% intersection over union (mAP@50), mean average precision from 50% to 95% intersection over union (mAP@50-95), and the inference time per image. Therefore, this work compares different model performance using these six evaluation metrics.
4.2.1 Params
in the model are determined by the architecture layer count, the number of neurons per layer, and the overall complexity of the model. In general, an increase in Params corresponds to a larger model size. While larger models generally offer better model performance, they also require more computational resources. Therefore, it is important to balance the model size with computational cost in practical applications.
4.2.2 FLOPs
representing floating-point operations, are commonly used to evaluate the computational complexity of the model. This metric serves as an indicator of the computational performance and speed of the models. In general, the models with lower FLOPs are more suitable for resource-limited environments, while the models with higher FLOPs require more powerful hardware support.
4.2.3 F1-Score
is the harmonic mean of the Precision and Recall of the model, ranging from 0 to 1. The value closer to 1 indicates that the model has achieved a better balance between Precision and Recall. Conversely, if either Precision or Recall is close to 0, F1-Score will also approach 0, signaling poor model performance.
4.2.4 mAP
is a widely used metric for evaluating the performance of object detection models. In object detection tasks, the model aims to recognize objects within an image and determines their locations. For each class, the model calculates the area under the Precision-Recall curve, known as average precision (AP), which measures the model detection performance for that specific class. The mAP score is calculated by averaging the AP values across all classes, providing an overall measure of the model detection accuracy.
4.3 Data Preparation
To ensure fairness in model performance comparison, the same data is used for all model training. We randomly divide the GRAZPEDWRI-DX [33] dataset into training, validation, and test sets in the ratio of 70%, 20%, and 10%, respectively. Specifically, the training set contains 14,204 images (69.88%); the validation set contains 4,094 images (20.14%), and the test set contains 2,029 images (9.98%).
Furthermore, data augmentation is performed on the training set before all models training. Specifically, this work adjusts the contrast and brightness of the X-ray images using the addWeighted function in the Open Source Computer Vision Library (OpenCV), and expanding the original training dataset from 14,204 images to 28,408 images.
4.4 Training
To ensure a fair comparison of the performance of different models on the GRAZPEDWRI-DX [33] dataset for pediatric wrist fracture detection, the same training environment and hardware (NVIDIA GeForce RTX 3090 GPUs) are used for all models training and evaluation. This work employs Python 3.9 and the PyTorch framework for training all of the models. For the hyperparameters of the model training, we set the batch size to 16 and the number of epochs of 100. In addition, all models are trained using the SGD optimizer, with the weight decay of 0.0005, momentum of 0.937, and the initial learning rate of 0.01.
4.5 Selection of Improved Methods
This work presents three different methods for improving YOLOv8, resulting in FCE-YOLOv8 models. Given four proposed FCE modules (i.e., SE, GC, GE, and GCT), we conduct the experiment to determine the most suitable improved method for each module. The experimental results are shown in Tables 123, and 4. As described in Section 3.1, we apply three methods (i.e., M1, M2, and M3) to train FCE-YOLOv8 with SE, GC, GE, and GCT modules, respectively. Apart from the “Method”, “Model Size”, and “Input Size” parameters, all other settings remain the same for all models.
Table 1 presents the experimental results for YOLOv8+SE. It can be seen that for the SE module, when the input image size is 640, choosing M2 to construct the YOLOv8+SE model can get the highest mAP@50 value of 64.08%; while the input image size is 1024, choosing M3 to construct the YOLOv8+SE model can get the highest mAP@50 value of 67.07%. In addition, with the input size of 640, Tables 23, and 4 show that the most suitable method for GC, GE, and GCT modules is all M1. YOLOv8+GC, YOLOv8+GE, and YOLOv8+GCT with M1 achieve the highest mAP@50 values of 64.55%, 64.15%, and 63.70%, respectively. However, when the input size is set to 1024, the most effective method for YOLOv8+GC is M3, resulting in the mAP@50 value of 66.32%. For both YOLOv8+GE and YOLOv8+GCT, M2 is the most effective, obtaining the mAP@50 values of 65.99% and 65.67%, respectively. Since the input image size is set to 1024 for comparisons with the SOTA models, as shown in Section 4.6, we select M3, M3, M2, and M2 for SE, GC, GE, and GCT, respectively, to construct FCE-YOLOv8.
4.6 Comparison with SOTA Models
For a comparative analysis with other SOTA models, this work sets the input image size to 1024 and configures all YOLOv8 models to large (L) and YOLOv9 to extended (E). The experimental results in Table 5 demonstrate that the proposed FCE-YOLOv8 models achieve satisfactory performance. Specifically, our proposed YOLOv8+GC-M3 and YOLOv8+SE-M3 models achieve the same F1-Score as YOLOv9, which is 0.66, reaching the SOTA level. In addition, the mAP@50 value of our proposed YOLOv8+SE-M3 model obtains the highest value of 67.07%, surpassing 65.78% of the previous SOTA model, YOLOv8+ResCBAM. Furthermore, our YOLOv8+GC-M3 model achieves the second-highest mAP@50 value of 66.32%. Based on this, the inference time of our YOLOv8+GC-M3 model is significantly shorter at 7.9 ms, compared to 8.7 ms of YOLOv8+ResCBAM and 16.1 ms of YOLOv9. For the model parameters and FLOPs, our proposed FCE-YOLOv8 models are also more efficient than the previous SOTA model. These experimental results demonstrate the positive effect of adding FCE modules to YOLOv8 on the model performance.
Fig. 6 presents the Precision-Recall Curve of our proposed FCE-YOLOv8 and other SOTA models for each class. According to Fig. 6, all models have good abilities to correctly detect the “fracture”, “metal”, and “text” classes, with the accuracy all above 90%. However, the poor ability of these models to detect the “bone anomaly” class seriously affects the mAP@50 values of the models. The proposed FCE-YOLOv8 models detect the “bone anomaly” class with the highest accuracy of 16.1%, compared to 15.0% of YOLOv8+ResCBAM, and 11.3% of YOLOv9. Details of this issue and its solution will be presented in Section 5.
In a real-world diagnostic scenario, to evaluate the accuracy improvement of YOLOv8 models enhanced with FCE modules for pediatric wrist fracture detection, four X-ray images are randomly selected from the used dataset. Fig. 7 presents the predictions of our proposed models and YOLOv8+ResCBAM, as well as the Ground-Truth images. Compared to the previous SOTA model (i.e., YOLOv8+ResCBAM), our proposed YOLOv8+GC and YOLOv8+SE models demonstrated greater accuracy in detecting “bone anomalies”, particularly in the first X-ray image on the left side of Fig. 7. This highlights the suitability of our proposed models as CAD tools to assist doctors and experts.
5 Discussion
Although the highest mAP@50 value of the proposed FCE-YOLOv8 models achieve 67.07% (surpassing the SOTA level), it falls short of the 70%. This limitation is primarily due to the poor performance of the model in predicting the “bone anomaly” and “soft tissue” classes. The GRAZPEDWRI-DX [33] dataset suffers from the class imbalance, for example, compared to 23,722 and 18,090 labels of the “text” and “fracture” classes, there are only 276 and 464 labels of the “bone anomaly” and “soft tissue” classes. The performance of YOLO models mainly depends on the quality and diversity of the training data. The limited training samples for these underrepresented classes result in the poor prediction ability of the model. Therefore, to further improve the model performance, obtaining more data for the “bone anomaly” and “soft tissue” classes is important. Our research team continues to search for additional data, and given the scarcity of pediatric-related X-ray images, we will immediately apply any newly available datasets to our FCE-YOLOv8 model training.
6 Conclusion and Future Work
Applying YOLOv8 [21], YOLOv8-AM [5, 23], and YOLOv9 [4] to pediatric wrist fracture detection has led to significant performance improvements on the publicly available GRAZPEDWRI-DX dataset [33]. On this dataset, the current SOTA model, YOLOv8+ResCBAM-L, obtains the mAP@50 values of 65.78%. However, this model is significantly larger than the original YOLOv8 model in terms of parameters and FLOPs. For example, the original YOLOv8-L model has only 43.61M parameters, while YOLOv8+ResCBAM-L has 53.87M, which is not satisfactory. Therefore, this paper proposes FCE-YOLOv8-L to improve the model performance effectively in a lightweight way. The parameters of all the proposed models are only from 43.64M to 43.85M, while the highest mAP@50 value reaches 67.07%, which exceeds the SOTA level.
To make this work available as CAD tools for medical image diagnostics, we plan to deploy our proposed FCE-YOLOv8 models as both web application and mobile phone application (e.g., Android and iOS) in the future, ensuring easy accessibility for doctors and experts.
7 Declarations
7.1 Author contributions statement
R.J. conceived and conducted the experiments, and wrote the main manuscript; C.C. contributed expertise on fractures and assisted in analyzing the results; E.X. performed the formal analysis and visualization; J.C. was responsible for funding acquisition, supervision, and manuscript editing. All authors have reviewed the manuscript.
7.2 Funding
This work is supported by National Science and Technology Council of Taiwan, under Grant Number: NSTC 112-2221-E-032-037-MY2.
7.3 Competing interests
The authors have no financial or proprietary interests in any material discussed in this article.
7.4 Ethics approval
This research does not involve human participants and/or animals.